# EarthMind: Towards Multi-Granular and Multi-Sensor Earth Observation with Large Multimodal Models

### v1

@@ 3.4 Training
EarthMind is trained via instruction tuning across diverse supervision formats, including VQA and segmentation tasks. The model learns to generate the target response conditioned on the image tokens, segmentation tokens, and the task instruction. Formally, the generation probability of the next token is defined as:
$\Pr\left(t_{i+1}\mid\underbrace{\texttt{<img>}_{1},\dots,\texttt{<img>}_{P}}_{\mathclap{\scriptstyle\text{image tokens}}},\ \underbrace{\texttt{<seg>}_{1},\dots,\texttt{<seg>}_{Q}}_{\mathclap{\scriptstyle\text{segmentation tokens}}},\ \underbrace{s_{1},\dots,s_{M}}_{\scriptstyle\text{instruction}},\ \underbrace{t_{1},\dots,t_{i}}_{\scriptstyle\text{generated response}};\ \boldsymbol{\Theta}\right)$ (5)
where $\boldsymbol{\Theta}$ denotes the trainable parameters of the LMM. For VQA tasks, we apply standard autoregressive training, minimizing the token-level cross-entropy loss over the ground-truth responses. For segmentation tasks, we use a combination of pixel-wise cross-entropy Loss and Dice Loss. In addition, the KL loss described in Sec. 3.2 is applied to supervise attention-based spatial priors.
Figure 3: Examples of EarthMind-Bench. There are ten tasks to evaluate the multi-sensor fusion ability of LMMs, covering perception and reasoning levels. The LLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on RGB, SAR or RGB-SAR fusion settings.
@@ 4 EarthMind-Bench
We observe that existing EO benchmarks lack support for multi-granular understanding and multi-sensor fusion. To address this limitation, we propose EarthMind-Bench, a new benchmark constructed from high-quality RGB–SAR paired data across various public datasets, including OpenEarthMap-SAR [44], DFC2023 Track2 [45], WHU-OPT-SAR [46], MSAW [47], and MultiResSAR [48]. We curate 2,000 samples from their test sets, and design a suite of 10 tasks spanning perception and reasoning, enabling a comprehensive evaluation of LMMs in EO scenarios, as shown in Fig. 3.
Perception. We include six tasks to evaluate both coarse- and fine-grained visual understanding, such as scene classification, object existence, object counting, and image captioning. In particular, we highlight two challenging tasks: 1) Hallucination detection, which requires models to answer some leading questions, and 2) Referring expression segmentation, where models must produce binary masks conditioned on natural language queries.
Reasoning. We organize reasoning tasks into two categories. First, spatial reasoning tasks include spatial relationship, where models infer the relative positions of given objects, and route planning, which requires generating a feasible path from a specified start point to an end point. Second, causal reasoning tasks consist of disaster forecasting and urban development assessment, where models are expected not only to make predictions or judgments, but also to provide reasoning evidence.
@@
All tasks are formatted as either multiple-choice (MC) or open-ended (OE) questions. For MC tasks, we report average accuracy across tasks as the evaluation metric. For OE tasks, we adopt GPT-based scoring to assess the quality of generated responses. All annotations are produced by human experts and have undergone a rigorous verification process, which is described in detail in Appendix D.
5 Experiments
5.1 Implementation Details
EarthMind builds upon InternVL2 [49] and SAM2 [28], adopting a curriculum learning strategy to fine-tune the model across three progressive stages. First, we enhance the instruction-following capability using 1.7M general image-text data, which covers image-level captioning, VQA, region-level object understanding and text-driven segmentation. Second, we introduce 1M EO-specific multimodal data to adapt EarthMind to the remote sensing domain. Third, we utilize our synthesized multi-sensor conversation corpus, and selectively retain examples from earlier stages to mitigate catastrophic forgetting. We fine-tune EarthMind with a learning rate of 4e-5 and a batch size of 2, training only the vision-language projector, the LLM via LoRA [50], and the mask decoder. All experiments are conducted on 8 NVIDIA A100-80G GPUs. More details about training datasets and details can be seen in Appendix F.
5.2 Benchmarks
In addition to our proposed EarthMind-Bench, we further evaluate the effectiveness of EarthMind on several widely-used EO benchmarks, assessing its capability in both multi-granularity and multi-sensor understanding. First, we conduct classification and vision-language evaluations on AID [51], UCMerced [52], RSVQA-HRBEN [53] and the VQA task of VRSBench [35] to perform image-level evaluation. AID is a large-scale aerial image classification dataset consisting of 10,000 images from 30 scene categories, collected via Google Earth. We follow the evaluation protocol of [14], using 20% of the dataset for testing. UCMerced contains 2,100 aerial images from 21 land use classes, commonly used for remote sensing scene classification. RSVQA-HRBEN consists of 47k question-answer pairs, in which we follow [14] to test on the “presence” and “Comparison” categories. The VQA part of VRSBench contains 123k samples covering diverse EO scenarios. Second, for region-level evaluation, we evaluate region-level understanding across region-level captioning and visual grounding task, where DIOR-RSVG [54] and visual grounding task of VRSBench are employed. Third, we conduct referring expression segmentation on the test sets of RefSegRS [55] and RRSIS-D [56] for pixel-level grounding, which include 1,817 and 3,481 test instances, respectively. Last, we involve the test set of BigEarthNet [57], SoSAT-LCZ42 [58], which consists of multispectral data and SAR ship detection dataset [59] to prove our multi-sensor understanding ability.
[…]
@@ 5.2 Benchmarks
Model Size M-Avg O-Avg Scene Class. Object Exist. Halluci. Detect. Object Count. Spatial Relation. Referring Segmen. Image Caption Disaster Forecast. Route Plann. Urban Assess.
@@
GeoChat [14] 7B 41.6 1.78 71.3 51.8 46.8 18.9 19.0 - 1.92 1.73 1.33 2.14
LHRS-bot [15] 7B 47.4 1.84 76.0 58.3 58.3 25.2 19.4 - 2.56 1.75 1.55 1.50
Skysensegpt [37] 7B 45.7 1.45 77.1 56.8 56.7 26.8 11.1 - 1.68 1.40 1.18 1.55
GeoPixel [32] 7B 59.4 2.06 87.2 67.8 73.6 33.5 34.7 46.8 2.80 1.80 1.68 1.95
EarthMind 4B 69.0 2.82 96.5 81.2 81.6 47.3 38.4 54.0 3.35 3.37 2.01 2.55
@@
GeoChat [14] 7B 40.2 1.45 58.1 49.8 46.8 27.9 18.5 - 1.78 1.65 1.25 1.56
LHRS-bot [15] 7B 41.0 1.71 58.5 56.3 48.5 23.5 18.1 - 1.86 1.70 1.45 1.83
Skysensegpt [37] 7B 39.9 1.55 51.9 53.8 49.2 33.8 10.7 - 1.76 1.70 1.35 1.38
GeoPixel [32] 7B 53.0 1.80 76.8 59.0 65.7 30.5 32.8 35.9 2.08 1.97 1.45 1.68
EarthMind 4B 67.5 2.64 95.4 77.4 74.6 46.8 43.1 53.0 3.10 3.25 1.89 2.30
Evaluation on RGB-SAR Fused modality
@@
EarthMind 4B 70.6 3.02 97.7 82.4 85.4 47.3 40.3 54.5 3.80 3.37 2.21 2.70
5.3 Main Results
Results on EarthMind-Bench. EarthMind-Bench supports evaluation under three settings: RGB only, SAR only, and RGB–SAR fusion. We compare EarthMind with state-of-the-art EO-specific LMMs and proprietary models such as GPT-4V [1] and GPT-4o [60]. Tab. 2 summarizes the results, from which we highlight three key findings: 1) Fine-grained and open-ended tasks remain challenging for existing LMMs. While coarse-grained tasks like scene classification are largely solved, tasks such as object counting and spatial relationship understanding still exhibit significant performance gaps. In particular, referring segmentation proves difficult for most models due to the lack of pixel-level reasoning ability. 2) Most models generalize poorly to SAR inputs. Performance under SAR-only settings lags behind RGB settings for nearly all models, likely due to training data limitations. In contrast, EarthMind demonstrates strong generalization to SAR inputs, benefiting from multi-sensor training. 3) Effective fusion requires learning cross-modal complementarity, not just stacking modalities. As the only open-source model with explicit fusion capability, EarthMind is compared against GPT-4 variants by feeding both RGB and SAR as pseudo-RGB inputs using the official multi-image interface. Results show that GPT-4 models experience performance degradation compared to RGB-only inputs, particularly on fine-grained tasks like Route Planning, Object Counting, and Spatial Relationship Understanding, where SAR data provides robust structural cues under adverse conditions. In contrast, EarthMind effectively captures cross-modal complementarity, yielding consistent improvements over single-modality inputs.
@@ 5.3 Main Results
Method Image-level Region-level AID UC RSVQA VRS-VQA RSVG VRS-VG GPT-4o [60] 74.7 88.8 - - - - LLaVA-1.5 [6] 72.0 84.4 63.1 76.4 - 5.1 GeoChat [14] 72.0 84.4 72.3 76.0 30.9 49.8 EarthGPT [17] - - 72.0 - 232.8 - EarthMarker [30] 78.0 86.5 - - 379.3 - EarthDial [31] 88.8 92.4 72.5 - - - EarthMind 97.2 95.0 74.0 78.9 428.2 55.6
@@
Method Pixel-level RRSIS-D RefSegRS LAVT [61] 56.8 47.4 RIS-DMMI [62] 60.3 52.2 Caris [63] 62.1 42.7 RM-SIN [56] 64.2 42.6 CroBIM [64] 64.5 59.8 GeoPixel [32] 67.3 - EarthMind 82.2 62.6
[…]
@@
Method Multispectral BigEarthNet SoSAT-LCZ42 GPT-4o [60] 49.0 15.5 Qwen2.5-VL†[65] 36.2 18.9 EarthDial [31] 69.9 60.7 EarthMind 70.4 58.3
@@
Method SAR Small Medium Large Single Multiple GPT-4o [60] 0.70 0.90 3.20 1.20 0 GeoChat†[14] 2.61 4.92 6.95 2.58 1.40 EarthDial [31] 12.14 26.02 35.56 26.03 6.06 EarthMind 13.58 28.55 36.78 27.45 6.99
@@
Results on Public Benchmarks. We evaluate EarthMind on a series of mainstream EO benchmarks (see Tab. 3). EarthMind consistently delivers strong performance across multiple levels of understanding. On image-level tasks such as scene classification (AID, UC-Merced) and EO visual question answering (RSVQA-HRBEN and VRSBench-VQA), it significantly outperforms previous models, including GPT-4o, despite using only 4B parameters. Moving to region-level tasks, EarthMind achieves a CIDEr score of 428.2 on DIOR-RSVG and 55.6% accuracy on VRSBench visual grounding, outperforming visual-prompt-based methods and highlighting the strength of our region encoder. When it comes to pixel-level benchmarks, EarthMind achieves top results on both RRSIS-D and RefSegRS, surpassing even specialized segmentation models and EO-focused LMMs. Beyond the RGB domain, EarthMind also demonstrates competitive results on multi-sensor understanding, as shown in Tab. 4, demonstrating the strong generalization towards complex EO scenarios.
Figure 4: (Left) Visualization of attention maps between segmentation tokens and image tokens within LLM. The top row is the results of “w/o SAP” and the bottom row is the results of “w/ SAP”, where the attention is reallocated to the region specified by the query. (Right) Comparison of our modality-mutual attention with other baselines under the “Token drop” study.
Table 5: Ablation on Spatial Attention Promoting (Left) and Cross-Modal Fusion (Right)
. Settings RefSegRS RRSIS-D mIoU $\uparrow$ oIoU $\uparrow$ mIoU $\uparrow$ oIoU $\uparrow$ w/o SAP 46.7 86.0 67.5 94.8 w/ SAP (first layer) 46.4 86.0 67.7 95.0 w/ SAP (middle layer) 48.0 86.9 71.8 95.5 w/ SAP (last layer) 47.2 86.5 69.2 95.0 w/ SAP (all layers) 48.4 87.4 72.0 96.0 Settings M-Avg $\uparrow$ mIoU $\uparrow$ w/o Modality Alignment 69.0 53.6 Naive Concatenate 68.4 52.7 Naive Attention 68.6 52.5 Ours 70.6 54.5

### v2

@@ 3.3 Training
Dataset Curation. Although large-scale visual instruction tuning datasets exist for single-modality remote sensing imagery, paired optical–SAR data with rich text annotations remain scarce. To address this gap, we construct a diversified set of QA pairs FusionEO to facilitate effective training. Specifically, we design a three-stage pipeline: (1) Metadata preparation. We leverage dataset-specific metadata (e.g., modality, category, geographic location, and relevant attributes) to provide essential contextual information. These metadata are then converted into concise captions through rule-based processing. (2) RoI-based summarization. To enrich the textual context with visual grounding cues, we incorporate region-level information from segmentation annotations. Foreground objects or localized regions of interest (RoIs) are explicitly highlighted in the images, thereby reducing hallucination risks. Mask-rendered images, together with the short captions from Stage 1, are then used as inputs to GPT-4o to generate more detailed and descriptive captions. (3) Self-instruct VQA generation. Finally, we extend the caption data into diverse VQA samples. GPT-4o is prompted in a few-shot manner, using seed examples to guide the generation of varied and semantically rich question–answer pairs. Additional implementation details are provided in Appendix G.
Training Objective. EarthMind is trained via instruction tuning across diverse supervision formats, including VQA and segmentation tasks. For VQA tasks, we apply standard supervised finetuning way, minimizing the token-level cross-entropy loss over the ground-truth responses. For segmentation tasks, we use a combination of pixel-wise cross-entropy Loss and Dice Loss.
Figure 3: Examples of EarthMind-Bench. There are ten tasks to evaluate the multi-sensor fusion ability of LMMs, covering perception and reasoning levels. The LLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on optical, SAR or optical-SAR fusion settings.
@@ 4 EarthMind-Bench
We observe that existing EO benchmarks cannot evaluate the performance of MLLMs under multi-sensor input. To address this limitation, we propose EarthMind-Bench, a new benchmark constructed from high-quality paired data with optical (including both RGB and multispectral) and SAR images across various public datasets, including BigEarthNet-MM (Sumbul et al., 2021), OpenEarthMap-SAR (Xia et al., 2025), DFC2023 Track2 (Persello et al., 2023), WHU-OPT-SAR (Li et al., 2022), MSAW (Shermeyer et al., 2020), and MultiResSAR (Zhang et al., 2025b). We curate 2,841 samples from their test sets, and design a suite of 10 tasks spanning perception and reasoning, enabling a comprehensive evaluation of MLLMs in EO scenarios, as shown in Fig. 3.
To comprehensively evaluate MLLMs in EO scenarios, we design a suite of perception and reasoning tasks. The perception tasks assess fundamental understanding, ranging from coarse-grained problems such as scene classification, object existence detection, and hallucination detection to more challenging tasks including object counting, image captioning, and referring expression segmentation. Building upon this foundation, the reasoning tasks are divided into spatial and causal categories. Spatial reasoning requires models to infer spatial relationships between objects or to perform route planning by generating a feasible path between a start and an end point. Causal reasoning, on the other hand, focuses on tasks such as disaster forecasting and urban development assessment, where models are expected not only to produce predictions or judgments but also to provide explanatory evidence supporting their reasoning.
@@
For text generation tasks, we report average accuracy on multiple-choice questions and adopt GPT-based scoring to evaluate the quality of open-ended responses. For mask generation tasks, segmentation quality is measured using mean Intersection-over-Union (mIoU). All annotations are produced by human experts and further subjected to a rigorous verification process, as described in Appendix E.
Table 1: Experimental results on EarthMind-Bench, in which different evaluation settings are employed to compare multi-sensor understanding ability. “M-Avg”: the average performance of multiple-choice tasks; “O-Avg”: the average performance of open-ended tasks. “Referring Segmen.” refers to the referring expression segmentation task, which is evaluated by mIoU. $\dagger$} denotes proprietary models. Bold means the best performance.
Model
Size
@@
M-Avg O-Avg Scene Class. Object Exist. Halluci. Detect. Object Count. Spatial Relation. Referring Segmen. Image Caption Disaster Forecast. Route Plann. Urban Assess.
@@
GeoChat (Kuckreja et al., 2024) 7B 35.5 1.78 40.9 51.8 46.8 18.9 19.0 - 1.92 1.73 1.33 2.14
LHRS-bot (Muhtar et al., 2024) 7B 41.3 1.84 45.5 58.3 58.3 25.2 19.4 - 2.56 1.75 1.55 1.50
Skysensegpt (Luo et al., 2024) 7B 39.8 1.45 47.4 56.8 56.7 26.8 11.1 - 1.68 1.40 1.18 1.55
GeoPixel (Shabbir et al., 2025) 7B 53.0 2.06 55.3 67.8 73.6 33.5 34.7 46.8 2.80 1.80 1.68 1.95
EarthDial (Soni et al., 2024) 4B 58.3 2.13 58.2 72.4 75.9 40.6 44.2 - 2.82 1.95 1.66 2.10
EarthMind 4B 61.7 2.82 64.3 77.5 83.6 50.1 33.1 56.0 3.35 3.37 2.01 2.55
@@
GeoChat (Kuckreja et al., 2024) 7B 34.3 1.45 28.6 49.8 46.8 27.9 18.5 - 1.78 1.65 1.25 1.56
LHRS-bot (Muhtar et al., 2024) 7B 35.1 1.71 28.9 56.3 48.5 23.5 18.1 - 1.86 1.70 1.45 1.83
Skysensegpt (Luo et al., 2024) 7B 34.2 1.55 23.5 53.8 49.2 33.8 10.7 - 1.76 1.70 1.35 1.38
GeoPixel (Shabbir et al., 2025) 7B 44.6 1.80 35.2 59.0 65.7 30.5 32.8 35.9 2.08 1.97 1.45 1.68
EarthDial (Soni et al., 2024) 4B 49.4 1.95 40.6 65.3 69.2 35.7 36.4 - 2.26 2.09 1.60 1.86
EarthMind 4B 61.3 2.64 64.4 77.4 74.6 46.8 43.1 53.0 3.10 3.25 1.89 2.30
Evaluation on Optical-SAR Fused modality
@@
EarthMind 4B 70.0 3.02 65.5 84.4 88.1 52.4 59.7 59.8 3.80 3.37 2.21 2.70
5 Experiments
5.1 Implementation Details
EarthMind builds upon InternVL2 (Chen et al., 2024b) and SAM2 (Ravi et al., 2024), adopting a progressive learning way including three steps. First, we enhance the instruction-following capability using 1.7M general image-text data, which covers image-level captioning, VQA, region-level object understanding and text-driven segmentation. Second, we introduce 1M EO-specific multimodal data to adapt EarthMind to the remote sensing domain. Third, we utilize our synthesized multi-sensor conversation corpus, and selectively retain examples from earlier stages to mitigate catastrophic forgetting. We fine-tune EarthMind with a learning rate of 4e-5 and a batch size of 2, training only the vision-language projector, the LLM via LoRA (Hu et al., 2022), and the mask decoder. All experiments are conducted on 8 NVIDIA A100-80G GPUs. More details about training datasets and details can be seen in Appendix G.
5.2 Benchmarks
In addition to our proposed EarthMind-Bench, we evaluate EarthMind on several widely-used EO benchmarks to assess its capability across diversified multi-sensor tasks. We conduct evaluations at three granularities: (1) Image-level: classification on AID (Xia et al., 2017) and UCMerced (Yang & Newsam, 2010), and VQA on RSVQA-HRBEN (Lobry et al., 2020) and VRSBench (Li et al., 2024b); (2) Region-level: captioning and visual grounding on DIOR-RSVG (Zhan et al., 2023) and VRSBench; (3) Pixel-level: referring expression segmentation on RefSegRS (Yuan et al., 2024b) and RRSIS-D (Liu et al., 2024b). Additionally, we evaluate multi-sensor understanding on SAR (Wang et al., 2019), BigEarthNet (Sumbul et al., 2019), and SoSAT-LCZ42 (Zhu et al., 2019) datasets.
@@ 5.2 Benchmarks
Method Image-level Region-level AID UC RSVQA VRS-VQA RSVG VRS-VG GPT-4o (OpenAI, 2024) 74.7 88.8 - - - - LLaVA-1.5 (Liu et al., 2023a) 72.0 84.4 63.1 76.4 - 5.1 GeoChat (Kuckreja et al., 2024) 72.0 84.4 72.3 76.0 30.9 49.8 EarthGPT (Zhang et al., 2024b) - - 72.0 - 232.8 - EarthMarker (Zhang et al., 2024a) 78.0 86.5 - - 379.3 - EarthDial (Soni et al., 2024) 88.8 92.4 72.5 - - - EarthMind 97.2 95.0 74.0 78.9 428.2 55.6
@@
Method Pixel-level RRSIS-D RefSegRS LAVT (Yang et al., 2022) 56.8 47.4 RIS-DMMI (Hu et al., 2023) 60.3 52.2 Caris (Liu et al., 2023b) 62.1 42.7 RM-SIN (Liu et al., 2024b) 64.2 42.6 CroBIM (Dong et al., 2024) 64.5 59.8 GeoPixel (Shabbir et al., 2025) 67.3 - EarthMind 82.2 62.6
Table 3: Quantitative performance on multispectral and SAR benchmarks. Evaluation follows the protocol of (Soni et al., 2024). $\dagger$} indicates that the results on benchmarks were reproduced using their official weights.
@@
Method Multispectral BigEarthNet SoSAT-LCZ42 GPT-4o (OpenAI, 2024) 49.0 15.5 Qwen2.5-VL†(Bai et al., 2025) 36.2 18.9 EarthDial (Soni et al., 2024) 69.9 60.7 EarthMind 70.4 58.3
@@
Method SAR Small Medium Large Single Multiple GPT-4o (OpenAI, 2024) 0.70 0.90 3.20 1.20 0 GeoChat†(Kuckreja et al., 2024) 2.61 4.92 6.95 2.58 1.40 EarthDial (Soni et al., 2024) 12.14 26.02 35.56 26.03 6.06 EarthMind 13.58 28.55 36.78 27.45 6.99
5.3 Results
EarthMind-Bench. EarthMind-Bench supports evaluation under three settings: Optical only, SAR only, and Optical–SAR fusion. For multispectral images, baseline models are evaluated using only RGB channels due to their inability to process multi-channel inputs, while EarthMind can leverage the full spectral information. We compare EarthMind with state-of-the-art EO-specific MLLMs and proprietary models such as GPT-4V (OpenAI, 2023) and GPT-4o (OpenAI, 2024). Tab. 1 summarizes the results, from which we highlight three key findings: 1) Fine-grained and open-ended tasks remain challenging for existing MLLMs. While coarse-grained tasks like scene classification are largely solved, tasks such as object counting and spatial relationship understanding still exhibit significant performance gaps. In particular, referring expression segmentation proves difficult for most models due to the lack of pixel-level reasoning ability. 2) Most models generalize poorly to SAR inputs. Performance under SAR-only settings lags behind Optical settings for nearly all models, likely due to training data limitations. In contrast, EarthMind demonstrates strong generalization to SAR inputs, benefiting from multi-sensor training. 3) Effective fusion requires learning cross-modal complementarity, not just stacking modalities. As the only open-source model with explicit fusion capability, EarthMind is compared against GPT-4 variants which feed both Optical and SAR data as pseudo-RGB inputs using the official multi-image interface. Results show that GPT-4 models experience performance degradation compared to Optical-only inputs, particularly on fine-grained tasks like Route Planning, Object Counting, and Spatial Relationship Understanding, where SAR data provides robust structural cues under adverse conditions. In contrast, EarthMind effectively captures cross-modal complementarity, yielding consistent improvements over single-modality inputs.
@@ 5.3 Results
Public Benchmarks. We evaluate EarthMind on mainstream EO benchmarks (Tab. 2). EarthMind consistently delivers strong performance across multi-level understanding tasks. On image-level tasks (AID, UC-Merced, RSVQA-HRBEN, VRSBench-VQA), it significantly outperforms previous models, including GPT-4o, despite using only 4B parameters. For region-level tasks, EarthMind achieves 428.2 CIDEr on DIOR-RSVG and 55.6% accuracy on VRSBench visual grounding. On pixel-level benchmarks, it achieves top results on both RRSIS-D and RefSegRS, surpassing specialized segmentation models. Beyond RGB data, EarthMind demonstrates competitive results on SAR and multispectral benchmarks (Tab. 3), showing strong generalization to diverse EO scenarios.
Table 4: (Left) Comparison of our proposed HCA methods with naive fusion methods on EarthMind-Bench, including Scene Classification (SC), Object Existence (OE), Hallucination Detection (HD), Object Counting (OC), Spatial Relationship (SR) and Referring Expression Segmentation (RS). (Right) Modality Attention Gap (Optical - SAR) across layers. For concatenation, the gap is computed from MAS using Eq. 1. For HCA, the gap is derived from the learned fusion weights ($w^{o}_{i}$ and $w^{s}_{i}$) during cross-modal attention. Lower gaps indicate better modality balance.
Method SC OE HD OC SR RS Avg Single Modality 64.4 77.5 83.6 50.1 43.1 56.0 62.5 Concatenation 64.2 77.9 83.3 49.6 34.2 57.6 61.1 Element-wise Average 64.7 79.3 85.4 49.8 36.9 57.0 62.2 Native Attention 65.0 79.9 85.2 49.8 38.7 58.2 62.8 HCA 65.5 84.4 88.1 52.4 59.7 59.8 68.3